dgiron23:

A bank account has a beginning balance of $560.00. After 6 months, the balance in the account has increased to $572.60. What interest rate is being earned annually on this account?

justus:

The interest entered was $12.60 \[SI=\frac{ PRT }{ 100 }\] \[12.60=\frac{ 560*R* 6 }{ 100 * 12}\] Now you can solve for R by removing it on the right \[\frac{ 12.60*100*12 }{ 560*6 }=R\]

justus:

Calculate it and let me know what you get. :)

dgiron23:

4.5% Thank you!

justus:

Correct, no problem!